# SBP-YOLO:A Lightweight Real-Time Model for Detecting Speed Bumps and Potholes

### v1

@@ 3.1 Experimental settings
The experiments were conducted on Ubuntu 24.04 with an Intel® Xeon E5-2680 CPU (2.4 GHz) and an NVIDIA Tesla V100 GPU (16 GB). The software environment comprised PyTorch 1.13.1 and CUDA 11.7. The model was trained using the Adam optimizer with an initial learning rate of 0.001, batch size of 16, and 300 epochs. Input images were resized to $\times 640$ 640 to enhance convergence and computational efficiency. Early stopping was employed, terminating training if validation performance did not improve for 30 consecutive epochs.
@@ 3.3 Evaluation indicators
[…]
where $TP$, $FP$, and $FN$ denote true positives, false positives, and false negatives, respectively. $P$ and $R$ represent precision and recall, and $M$ is the number of object categories.
@@ 3.4 Ablation experiment
Model P (%) R (%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50\text{-}95}$ GFLOPs Params (M) FPS
@@ 3.5 Knowledge Distillation Experiment
In our study, a teacher model with similar architecture and S-level scale (20.6 GFLOPs, 7.42M Params) was employed to guide the training of the student model. As shown in Table 3, knowledge distillation leads to consistent improvements across all metrics. In particular, $\mathrm{mAP}_{50\text{-}95}$ increased by 3.5% on the validation set and 3.4% on the test set, indicating not only enhanced detection performance but also improved generalization. The comparable gains across both sets suggest that the student model benefits from the teacher’s guidance without exhibiting overfitting.
@@ 3.6 Model comparison experiment
To comprehensively evaluate detection quality and computational efficiency, we compare SBP-YOLO with representative YOLO variants (YOLOv5–YOLOv12) across standard performance and complexity metrics. As reported in Table 4, SBP-YOLO achieves an $\mathrm{mAP}_{50}$ of 86.6%, surpassing the YOLOv11n baseline by 5.4%, while reducing model parameters, FLOPs, and storage by 20.9%, 7.9%, and 17.9%, respectively. With knowledge distillation, SBP-YOLO (Distilled) further improves detection accuracy, reaching 87.0% $\mathrm{mAP}_{50}$ and 54.6% $\mathrm{mAP}_{50\text{-}95}$, without introducing additional computational overhead. These results highlight the effectiveness of our architecture and training strategy in achieving a strong trade-off between accuracy and efficiency.
@@ 4.1 Experimental Environment
Figure 9: Experimental Hardware Platform
@@ 5 Conclusion
This paper presents SBP-YOLO, a lightweight and efficient object detection framework designed for real-time detection of speed bumps and potholes to facilitate predictive control in intelligent suspension systems. Building upon the YOLOv11 architecture, SBP-YOLO introduces a P2 detection head to improve sensitivity to small and distant targets, integrates GhostConv and VoVGSCSPC modules for efficient multi-scale feature extraction, and adopts proposed LEDH to reduce computational overhead without compromising accuracy. To further enhance robustness and generalization in dynamic environments, a hybrid training strategy is employed, combining NWD loss, backbone-level knowledge distillation (BCKD), and Albumentations-based augmentation. This strategy yields consistent improvements in detection performance, particularly for small-scale anomalies, outperforming the YOLOv11n baseline by 5.8% in mAP. Experimental results show that, following TensorRT FP16 optimization, SBP-YOLO achieves 139.5 FPS on the Jetson AGX Xavier platform, representing a 12.4% increase in inference speed compared to the baseline YOLOv11n model with the same P2 detection head structure (124.1 FPS). This confirms the efficiency advantage of the presented architectural enhancements for embedded real-time deployment in intelligent suspension systems.

### v2

@@ 3.1 Experimental settings
The experiments were conducted on Ubuntu 24.04 with an Intel® Xeon E5-2680 CPU (2.4 GHz) and an NVIDIA Tesla V100 GPU (16 GB). The software environment comprised PyTorch 1.13.1 and CUDA 11.7. The model was trained using the Adam optimizer with an initial learning rate of 0.001, batch size of 16, and 300 epochs. Input images were resized to $640\times 640$ to enhance convergence and computational efficiency. Early stopping was employed, terminating training if validation performance did not improve for 30 consecutive epochs.
@@ 3.3 Evaluation indicators
$\mathrm{AP}=\int_{0}^{1}P(R)\,dR,\quad\mathrm{mAP}=\frac{1}{M}\sum_{i=1}^{M}\mathrm{AP}_{i}$ (12)
where $TP$, FPFP, and $FN$ denote true positives, false positives, and false negatives, respectively. $P$ and $R$ represent precision and recall, and $M$ is the number of object categories.
@@ 3.4 Ablation experiment
Model P (%) R (%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50\text{-}95}$ GFLOPs Params (M) FPSb=1
@@ 3.5 Knowledge Distillation Experiment
In our study, a teacher model with similar architecture and S-level scale (20.6 GFLOPs, 7.42M Params) was employed to guide the training of the student model. As shown in Table 3, KD leads to consistent improvements across all metrics. In particular, $\mathrm{mAP}_{50\text{-}95}$ increased by 3.5% on the validation set and 3.4% on the test set, indicating not only enhanced detection performance but also improved generalization. The comparable gains across both sets suggest that the student model benefits from the teacher’s guidance without exhibiting overfitting.
@@ 3.6 Model comparison experiment
To comprehensively evaluate detection quality and computational efficiency, we compare SBP-YOLO with representative YOLO variants (YOLOv5-YOLOv12) across standard performance and complexity metrics. As reported in Table 4, SBP-YOLO achieves an $\mathrm{mAP}_{50}$ of 86.6%, surpassing the YOLOv11n baseline by 5.4%, while reducing model parameters, FLOPs, and storage by 20.9%, 7.9%, and 17.9%, respectively. With knowledge distillation, SBP-YOLO (Distilled) further improves detection accuracy, reaching 87.0% $\mathrm{mAP}_{50}$ and 54.6% $\mathrm{mAP}_{50\text{-}95}$, without introducing additional computational overhead. These results highlight the effectiveness of our architecture and training strategy in achieving a strong trade-off between accuracy and efficiency.
@@ 4.1 Experimental Environment
Figure 9: Experimental hardware setup: (a) experimental platform; (b) display terminal for real-time inference results; (c) Jetson AGX Xavier-based hardware configuration .
@@ 5 Conclusion
This paper presented SBP-YOLO, a lightweight framework for real-time detection of speed bumps and potholes to support predictive suspension control. Building on YOLOv11, it incorporates a P2 detection head with LEDH employing shared GroupConv transformations to enhance small-object sensitivity while reducing computational overhead, alongside GhostConv and VoVGSCSPC modules for efficient multi-scale feature extraction. A hybrid training strategy combining NWD loss, backbone-level knowledge distillation, and advanced data augmentation improves robustness and generalization, achieving a 5.8% mAP gain over the YOLOv11n baseline. After TensorRT FP16 optimization, the model reaches 139.5 FPS on Jetson AGX Xavier, a 12.4% speedup over P2-enhanced YOLOv11n while maintaining superior accuracy. These results demonstrate the framework’s effectiveness for real-time embedded deployment, enabling anticipatory perception and adaptive damping control, with future work exploring infrared integration and multimodal detection to further enhance robustness under diverse conditions.